There were 510 people at a play. The admission price was $3.00 for adults and $1.00 for children. The admission receits were $930. How many adults and children attended.
Answer by mananth(16946)

Adult x Tickets
child y Tickets

1 x + 1 y = 510 .............1
3 x + 1 y = 930 .............2
Eliminate y
multiply (1)by -1
Multiply (2) by 1
-1 x -1 y = -510
3 x 1 y = 930
Add the two equations
2 x = 420
/ 2
x = 210
plug value of x in (1)
1 x + 1 y = 510
210 + 1 y = 510
1 y = 510 -210
1 y = 300
y = 300
Adult 210 Tickets
child 300 Tickets
m.ananth@hotmail.ca